# Bondal-Orlov’s reconstruction theorem in noncommutative projective geometry

### v1

@@ 4. Bondal-Orlov’s reconstruction theorem for noncommutative projective schemes
If the Gorenstein parameters of $A,B$ are $l_{A},l_{B}$ and the global dimensions of $A,B$ are $n_{A},n_{B}$, then $A^{[l_{A}]},B^{[l_{B}]}$ are AS-regular algebras of dimensions $n_{A},n_{B}$ over $\left(A^{[l_{A}]}\right)_{0},\left(B^{[l_{B}]}\right)_{0}$, respectively ([mori2021acategorical, Corollary 4.4]). In addition, the Gorenstein parameters of $A^{[l_{A}]},B^{[l_{B}]}$ are $1$. From 4.20, we have the balanced dualizing complexes of $A^{[l_{A}]},B^{[l_{B}]}$ given by $A^{[l_{A}]}_{\nu_{1}}(-1)[n_{A}],B^{[l_{B}]}_{\nu_{2}}(-1)[n_{B}]$ for some graded automorphisms $\nu_{1},\nu_{2}$ of $A^{[l_{A}]},B^{[l_{B}]}$, respectively. This means that $-\otimes\mathcal{K}_{A^{[l_{A}]}},-\otimes\mathcal{K}_{B^{[l_{B}]}}$ are AZ-anti-ample. Moreover, we have
@@
from 3.8. So, we apply 4.18 to obtain $\operatorname{qgr}\left(A^{[l_{A}]}\right)\simeq\operatorname{qgr}\left(B^{[l_{B}]}\right)$ and then $\operatorname{qgr}(A)\simeq\operatorname{qgr}(B)$.

### v2

@@ 4. Bondal-Orlov’s reconstruction theorem for noncommutative projective schemes
If the Gorenstein parameters of $A,B$ are $l_{A},l_{B}$ and the global dimensions of $A,B$ are $n_{A},n_{B}$, then $A^{[l_{A}]},B^{[l_{B}]}$ are AS-regular algebras of dimensions $n_{A},n_{B}$ over $\left(A^{[l_{A}]}\right)_{0},\left(B^{[l_{B}]}\right)_{0}$, respectively ([mori2021acategorical, Corollary 4.4]). In addition, the Gorenstein parameters of $A^{[l_{A}]},B^{[l_{B}]}$ are $1$. From 4.22, we have the balanced dualizing complexes of $A^{[l_{A}]},B^{[l_{B}]}$ given by $A^{[l_{A}]}_{\nu_{1}}(-1)[n_{A}],B^{[l_{B}]}_{\nu_{2}}(-1)[n_{B}]$ for some graded automorphisms $\nu_{1},\nu_{2}$ of $A^{[l_{A}]},B^{[l_{B}]}$, respectively. This means that $-\otimes\mathcal{K}_{A^{[l_{A}]}},-\otimes\mathcal{K}_{B^{[l_{B}]}}$ are AZ-anti-ample. Moreover, we have
@@
from 3.8. So, we apply 4.20 to obtain $\operatorname{qgr}\left(A^{[l_{A}]}\right)\simeq\operatorname{qgr}\left(B^{[l_{B}]}\right)$ and then $\operatorname{qgr}(A)\simeq\operatorname{qgr}(B)$.